two cars start from the same point and travel in different directions. the car traveling west leaves 1 hour (earlier/later) than the car traveling east. the eastbound car travels 40 mph, and the westbound travels 60 mph. When they are 240 miles apart, how long had they traveled?
Answer by richwmiller(17219)

earlier
40x+60(1+x)=240
40x+60+60x=240
100x+60=240
100x=180
x=1.8 hours at 40 and 2.8 hours at 60
later
40x+60(x-1)=240
40x-60+60x=240
100x-60=240
100x=300
x=3 hours at 40=120
x-1 =2 hours at 60=120